OpenStudy (anonymous):

determine when to use the rolles theorem f(x)=(x-1)(x-2)(x-3) [1,3]

OpenStudy (anonymous):

sshould i foil everythibng first?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you only need to check to see if \[f(1)=f(3)\]

OpenStudy (anonymous):

which is a lot easier to check in factored form, because you can see almost instantly that \[f(1)=f(3)=0\]

OpenStudy (anonymous):

i am not exactly sure what "when to use rolles theorem" means if it means "is rolles theorem applicable" then the answer would be "yes"

OpenStudy (anonymous):

because x cant be 1 or 3, thats why f(1)=f(3)?

OpenStudy (anonymous):

what if i were to find the derivative?

OpenStudy (anonymous):

should i use the product rule?